without calculating, solve for z and explain why
3x + 2y + 4z -7w = 4
2x -5y +7z + 2w = 7
4x +2y -3z -3w = -3
x -11y +2z +9w = 2

Answer by Alan3354(69443)

without calculating, solve for z and explain why
3x + 2y + 4z -7w = 4
2x -5y +7z + 2w = 7
4x +2y -3z -3w = -3
x -11y +2z +9w = 2
==========================
The coefficient of z equals the right side.
--> z = 1
x = y = w = 0